find x.
LMN is isosceles, L is the vertex angle, LM = 3x -2, LN =2x + 1, and MN = 5x -2

Answer by nycsub_teacher(90)

Isosceles triangles have two equal sides and one side that is not equal.
Side LM = side LN.
3x - 2x = 2x + 1
3x - 2x = 2 + 1
x = 3